OpenStudy (anonymous):

Given that f(x) = x2 - 3x + 3 and g(x) = , solve for f(g(x)) when x = 5

OpenStudy (jdoe0001):

f(x) = x2 - 3x + 3 and g(x) = , <--- \(\large ? \)

OpenStudy (anonymous):

\[\frac{ x-1 }{ 4}\]

OpenStudy (anonymous):

@jdoe0001

OpenStudy (ranga):

f(g(x)) when x = 5 First find g(x) when x = 5. Then put that value as x in f(x) and evaluate f(x)